A manufacturer of flashlight batteries claim that the average life of his product will exceed 40 hours. A company is willing to buy a very large shipment of batteries if the claim is true. A random sample of 36 batteries is tested, and it is found that the sample mean is 45 hours. If the population of the batteries has a standard deviation of 5 hours, is it likely that the batteries will be bought?
1
Expert's answer
2021-09-30T23:41:16-0400

"H_0: \mu ≤ 40 \\

H_1: \mu > 40 \\

\bar{x} = 45 \\

\sigma= 5 \\

n = 36"

Right-tailed test.

Test-statistic:

"Z = \frac{\bar{x} -\mu}{\sigma / \sqrt{n}} \\

Z = \frac{45-40}{5 / \sqrt{36}} = 6.00"

Let use 0.01 level of significance.

"P-value=P(Z>6.0) \\

= 1 -P(Z<6.0) \\

= 1-0.999968 \\

= 0.000032"

Let us use α=0.01

Reject H0 if p-value < α.

P-value = 0.000032 < α = 0.01

We reject H0.

We can conclude that the average life of his product will exceed 40 hours.

A company will buy a very large shipment of batteries.